OpenStudy (anonymous):

hAntonio is seven less than four times his daughter's age. Antonio is 45; how old is his daughter? 13 15 17 19

OpenStudy (wasiqss):

13

OpenStudy (anonymous):

" seven less than for times \(x\)" translates in to math as \[4x-7\] solve \[4x-7=45\] add 7 divide by 4

OpenStudy (anonymous):

you will in fact get 13 via \[4x-7=45\] \[4x=45+7=52\] \[x=52\div4=13\]